Good morning distinguished Tutors,i need help with this question-- (a) Determine the maximum and minimum points and sketch the graph of y=x(12-2x)^2
(b)Determine the area bounded by y=x(12-2x)^2,the x axis and the lines x=(-1) and x=5.


Answer by addingup(3677)

max{y= x(12-2x)^2}= 128 at x = 2
min{y = x(12-2x)^2}= 0 at x= 6
---------
Since you're dealing with a square you'll need two plots.